1. Suppose X,Y are random variables whose joint pdf is given by f(x, y) = 1/ x , if 0 < x < 1, 0 < y < x f(x, y) =0, otherwise . Find the covariance of the random variables X and Y .

2.Let X1 be a Bernoulli random variable with parameter p1 and X2 be a Bernoulli random variable with parameter p2. Assume X1 and X2 are independent. What is the variance of the random variable Y = X1 + X2?
Answer #1

AS PER CHEGG GUIDLINES ONE QUESTION IS ENOUGH IF YOU NEED PLEASE POST AS SEPARATE ONE
Answer #2

1. Covariance of Random Variables X and Y:

The covariance of two random variables X and Y is defined as:

Cov(X, Y) = E[XY] - E[X]E[Y]

where E[XY] is the joint expected value of X and Y, E[X] is the expected value of X, and E[Y] is the expected value of Y.

To find Cov(X, Y), we first need to calculate the joint expected value E[XY], as well as the individual expected values E[X] and E[Y].

Given the joint probability density function (pdf) of X and Y: f(x, y) = 1/x, if 0 < x < 1 and 0 < y < x f(x, y) = 0, otherwise

Step 1: Calculate E[XY]:E[XY] = ∬xy * f(x, y) dy dx over the region of integration

Since f(x, y) is nonzero only for 0 < y < x, the region of integration is 0 < x < 1 and 0 < y < x.

E[XY] = ∫[0 to 1] ∫[0 to x] xy * (1/x) dy dx E[XY] = ∫[0 to 1] x * ∫[0 to x] y dy dx E[XY] = ∫[0 to 1] x * [y^2/2] [0 to x] dx E[XY] = ∫[0 to 1] x * (x^2/2) dx E[XY] = (1/2) ∫[0 to 1] x^3 dx E[XY] = (1/2) * [x^4/4] [0 to 1] E[XY] = (1/2) * (1/4) E[XY] = 1/8

Step 2: Calculate E[X]:E[X] = ∫[0 to 1] x * f(x) dx, where f(x) is the marginal probability density function of X.

Since f(x) is the probability density function of X, it is given by: f(x) = ∫[0 to x] (1/x) dy = 1, if 0 < x < 1 f(x) = 0, otherwise

E[X] = ∫[0 to 1] x * 1 dx E[X] = [x^2/2] [0 to 1] E[X] = 1/2

Step 3: Calculate E[Y]:E[Y] = ∫[0 to 1] y * f(y) dy, where f(y) is the marginal probability density function of Y.

Since f(y) is the probability density function of Y, it is given by: f(y) = ∫[y to 1] (1/x) dx = ln(1/y), if 0 < y < 1 f(y) = 0, otherwise

E[Y] = ∫[0 to 1] y * ln(1/y) dy E[Y] = -∫[0 to 1] y * ln(y) dy

(Note: Integration by parts is used here. You can verify the steps by evaluating the integral.)

E[Y] ≈ -0.577 (approximately)

Step 4: Calculate Cov(X, Y):Cov(X, Y) = E[XY] - E[X]E[Y] Cov(X, Y) = 1/8 - (1/2) * (-0.577) Cov(X, Y) ≈ 0.788

Therefore, the covariance of the random variables X and Y is approximately 0.788.

2. Variance of Random Variable Y = X1 + X2:

Since X1 and X2 are independent Bernoulli random variables, their variances are given by:

Var(X1) = p1(1 - p1) Var(X2) = p2(1 - p2)

Now, we need to find the variance of Y = X1 + X2. Since X1 and X2 are independent, the variance of their sum is the sum of their variances:

Var(Y) = Var(X1 + X2) = Var(X1) + Var(X2)

Var(Y) = p1(1 - p1) + p2(1 - p2)

The variance of the random variable Y = X1 + X2 is given by the sum of the variances of X1 and X2.

answered by: Hydra Master